OpenStudy (anonymous):

polar to cartesian and cartesian to polar. can someone explain this to me? need your help for my incoming exam

OpenStudy (chaise):

Polar -> Cart x=rcos(theta) y=rsin(theta) Cart -> Polar x^2+y^2=r^2 theta=arctan(y/x)

OpenStudy (anonymous):

for example 4x^2 + 9y^2 = 36.. the answer was r^2 = 36/ 4cos^2 theta + 9sin^2 theta. what is the process?

OpenStudy (anonymous):

yes yes substituting but i always stuck up..

OpenStudy (anonymous):

So you're trying to find tricks that will prove worthy when you come across these type of equations?

OpenStudy (anonymous):

i don't know. ive always tried to convert that equation without seeing the real answer but i always end up nothing..

OpenStudy (anonymous):

Cartesian from my experience are basically point values in a 3 D graph like xa +yb + zk. that pinpoints where it's at in an environment. While polar coordinates resemble more for objects that you want to find by circular means like finding a pin point on the surface of the earth if say you were in the middle of the core where the origin of the graph should be. with r = telling you far from the earth's core is the surface point and theta being at which angle from the horizontal axis

OpenStudy (anonymous):

yes yes but what about 4x^2 + 9y^2 = 36.. converting to polar

OpenStudy (anonymous):

Draw the circle everytime! :)  Then you just have to remember good ol' Pythagoras and trig, since r^2 is just the hypotenuse. r^2 = x^2 + y^2

OpenStudy (anonymous):

In that instance just see that 4x^2 + 9y^2 = 36 fits into (2x)^2+(3y)^2=6^2 Your r^2 would be the radius of a circle that contains the points cartesian points (X,Y), where X = 2x and Y = 3y

OpenStudy (anonymous):

so then \[\tan \theta = \frac{ X }{ Y } = \frac{ 2x }{ 3y }\]

OpenStudy (anonymous):

derp, y/x

OpenStudy (anonymous):

\[\theta = \arctan \frac{ Y }{ X } = \arctan \frac{ 3y }{ 2x }\] and 2x = r sin theta 3y = r sin theta

OpenStudy (anonymous):

okay then.. how come the answer is r^2 = 36/ 4cos^2 theta + 9sin^2 theta

OpenStudy (anonymous):

is it 36 / (4cos^2theta + 9sin^2theta)? With those parenthesis there?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's because I'm dumb and like circles more than sense :P With 4x^2 + 9y^2 = 36, like you said we're trying to find r. I thought that we could just make up a new R, but that doesn't help in this case. The only things we know for sure about the equation are our original polar definitions x = r cos theta y = r sin theta So we can plug that into that equation directly 4(r cos theta)^2 + 9(r sin theta)^2 = 36 Then you can pull out the r's from those terms and get r^2 (4cos^2 theta + 9 sin^2 theta) = 36 and that's where they got that answer. r^2 = 36/ (4cos^2 theta + 9 sin^2 theta) Hope I didn't confuse you before :/

OpenStudy (anonymous):

okay i didnt pulled out r and everything went wrong. OMG thanks!!!

OpenStudy (anonymous):

:) Welcome